# HiT: Hierarchical Transformers for Unsupervised 3D Shape Abstraction

### v1

@@ 1 Introduction
It is well established in cognitive science that humans perceive shapes as structured collections of parts, organized hierarchically [29, 16]. This hierarchical part cognition helps to reason about the function of each part [24], and allows humans to effortlessly establish correspondences between similar parts in different shapes in a collection. Introducing similar hierarchical decompositions to 3D digital shapes not only aligns representations with human perception, but also enables key applications in computer graphics and robotics. In graphics, it facilitates part-aware editing [14], attribute transfer [42], and compositional shape generation [3]; in robotics, it can support affordance reasoning, manipulation planning [21], and generalizable interaction with objects [34], in a coarse-to-fine manner.
@@ 1.1 Structured neural shape representations
A common approach to 3D shape understanding is to reconstruct them as compositions of primitives or semantic parts. Prior works learn such part-aware representations by approximating implicit surfaces with fixed or learned primitives. For example, [30, 37, 12, 13] represent shapes as unions of superquadrics or Gaussians, but their fixed structures often fail to capture surfaces accurately. BAE-Net [5], Neural-Parts [31], and DAE-Net [7] improve flexibility using MLPs to represent parts, either through branching layers or by parameterizing deformations of simple shapes with invertible neural networks as learned homeomorphisms. While more expressive, these methods often yield coarse decompositions and can collapse without good initialization.
@@ 1.2 Shape segmentation with pre-trained models
Several recent methods infer 3D segmentation by importing information external to raw geometry. Multi‑view neural‑field approaches lift 2D semantic or panoptic masks from posed images into an implicit 3D volume, achieving scene‑level labels without part annotations [45, 40, 38, 19, 33, 17]. Complementary work leverages vision-language training to segment shapes under zero‑shot or open‑vocabulary setups, from point‑cloud labeling [23, 4, 18] to CLIP‑based mesh highlighting [11, 8, 1]. All of these methods require either 2D mask supervision or large language–vision models, whereas our method learns part hierarchies directly from geometry without any auxiliary cues.
@@ 2 Method
In section˜2.1, we describe how decoder layers use part codebooks and cross-attention to define part–subpart relationships. In section˜2.2, we show how parts are geometrically grounded using 3D convex primitives with nested containment. Finally, section˜2.3 outlines the training objectives that combine reconstruction loss with regularizers to achieve self-supervised training.
@@ 3 Experiments
We validate our method through the task of part segmentation (section˜3.1), where we show our method can outperform all previous baselines in recovering consistent parts in all categories. We then show how our hierarchical part reconstruction achieves better reconstruction metrics in finer levels while higher levels remain semantically meaningful level of shape abstraction (section˜3.2). Finally, we ablate our design choices and analyze our training dynamics and recovered codebooks (section˜3.3).
Implementation details
We train our model for 150 epochs using a batch size of $32$ and a learning rate of $10^{-4}$. All categories are trained jointly. Input point clouds are uniformly sub-sampled to $2048$ points, and the voxel output resolution of the encoder is set to $32$. All the meshes are extracted using marching cubes at a resolution of $128$ and a threshold of $0.5$. We train the model across $4$ hierarchy levels, with the number of parts per level set to $[4,8,16,32]$, unless specified otherwise. Each convex uses $32$ planes, while the latent code dimensionality is $64$. We set the following loss weights: $\lambda_{\text{contain}}{=}0.01$, $\lambda_{\text{balance}}{=}0.01$, and $\lambda_{\text{cvxnet}}{=}0.01$.
@@ 3.1 Part segmentation – figures˜3 and 4
We train our method on $55$ categories from the ShapeNet-v2 dataset, following the train-test split used by Zhang et al. [44]. Our approach supports a unified training scheme, allowing all categories to be trained jointly. We further analyze this design in section˜3.3, where we compare it to category-specific training. For segmentation accuracy evaluation, since the ShapeNet [2] dataset does not provide segmentation labels, we leverage the fine-grained part annotations from the PartNet [26] dataset. This differs from previous methods, which evaluate segmentation accuracy on the ShapeNetPart dataset [41], containing only coarse segmentation ground-truth labels. The Door, Scissors, and Refrigerator categories appear only in PartNet and not in ShapeNet; hence for these, we use all instances for evaluation only. Each ground-truth point cloud for evaluation in PartNet dataset contains $10,000$ points.
@@
Our method significantly outperforms all baselines in part segmentation, both quantitatively and qualitatively. It produces consistent labeling of corresponding parts across instances within a category, whereas baselines often miss parts entirely or misclassify semantically similar regions. Additionally, qualitative results on the Table and Chair categories highlight how our flexible tree structure adapts to varying geometries in a category, enabling different numbers of subparts as needed.
@@ 3.2 Hierarchical reconstruction – figure˜5
We demonstrate that our model captures coherent high-level abstractions at early levels and progressively refines fine-grained details at deeper levels. Following the protocol of [44], we train and test on ShapeNet-v2 using the official splits. For evaluation, we compare reconstructed meshes—sampled to 100,000100{,}000 points each—against ground-truth point clouds, using symmetric Chamfer Distance and voxelized IoU at $128^{3}$ resolution. Our hierarchy employs $[6,16,24,36]$ parts. We compare against BAENet [5], RIMNet [28], and DAENet [7]. While BAENet and DAENet are single-level methods, RIMNet supports hierarchical reconstruction but only via a fixed binary tree, which yields coarse abstractions and loss of detail in later levels (e.g., rifles in figure˜5). By contrast, our model recovers flexible multi-branch hierarchies that balance abstraction and detail across levels.
3.3 Ablation study – tables˜1, 7 and 6
@@
We ablate our design choices by analyzing the effect of each loss, showing all loss components contribute to the quality of the decomposition. While balance and containment losses have little effect on quantitative metrics, removing them yields imbalanced and non–self-contained hierarchies, as seen in the qualitative results. Further, we analyze our training regime. While our method offers an effective, convenient, and generalized training on all object categories at once, we further analyze how a per-category training similar to previous works affects our results. This shows that a per-category training can result in an improved specialized model for each category, in exchange for more time and compute. Finally, we provide a t-SNE analysis of the learned part embeddings, demonstrating that embeddings of similarly labeled parts cluster closely together.
@@ 4 Conclusion
We introduce HiT, a self-supervised attention-based hierarchical neural field representation that learns general shape abstractions in a coarse-to-fine manner across diverse categories. Our core contribution is a novel cross-attention mechanism. This design enables the dynamic discovery of parent–child relationships across levels and enables learning tree-structured hierarchies. To the best of our knowledge, our method is the first to enable general cross-category hierarchical shape abstraction.
However, it is limited by the requirement that the number of parts at each hierarchy level must be fixed, which can occasionally cause unnatural decompositions. Furthermore, having a large number of convexes at finer levels can sometimes cause over-segmentation at the leaf nodes. A promising future direction is to make the number of convexes adaptive, for example by selecting them based on sparsity or reconstruction utility, so that the hierarchy remains compact and semantically meaningful. Another exciting direction is to extend HiT to generative modeling, where hierarchical structure can provide a powerful prior for shape synthesis. We hope this work serves as a step toward more general and interpretable 3D shape abstraction.
@@ Appendix A Supplementary
In the supplementary material, we present the hierarchical decomposition results of our method on the Objaverse dataset [9]. As shown in the examples below, our model dynamically allocates a different number of convexes depending on the input shape. Figures 8, 9, 10, 11, and 12 provide qualitative results of these decompositions. As can be seen, our model generates diverse decompositions tailored to the structural characteristics of each shape.

### v2

@@ 1 Introduction
It is well established in cognitive science that humans perceive shapes as structured collections of parts, organized hierarchically [31, 17]. This hierarchical part cognition helps to reason about the function of each part [25], and allows humans to effortlessly establish correspondences between similar parts in different shapes in a collection. Introducing similar hierarchical decompositions to 3D digital shapes not only aligns representations with human perception, but also enables key applications in computer graphics and robotics. In graphics, it facilitates part-aware editing [15], attribute transfer [48], compositional shape generation [3], and motion generation [45, 51]. In robotics, it can support affordance reasoning, manipulation planning [22], and generalizable interaction with objects [37], in a coarse-to-fine manner.
@@ 2.1 Structured neural shape representations
A common approach to 3D shape understanding is to reconstruct them as compositions of primitives or semantic parts. Prior works learn such part-aware representations by approximating implicit surfaces with fixed or learned primitives. For example, [32, 41, 13, 14, 44, 28] represent shapes as unions of superquadrics or Gaussians, but their fixed structures often fail to capture surfaces accurately. BAE-Net [7], Neural-Parts [33], DAE-Net [5] and PartSDF  [39] improve flexibility using MLPs to represent parts, either through branching layers or by parameterizing deformations of simple shapes with invertible neural networks as learned homeomorphisms. While more expressive, these methods often yield coarse decompositions and can collapse without good initialization.
@@ 2.2 Shape segmentation with pre-trained models
Several recent methods infer 3D segmentation by importing information external to raw geometry. Multi‑view neural‑field approaches lift 2D semantic or panoptic masks from posed images into an implicit 3D volume, achieving scene‑level labels without part annotations [52, 46, 42, 20, 36, 18, 12]. Complementary work leverages vision-language training to segment shapes under zero‑shot or open‑vocabulary setups, from point‑cloud labeling [24, 4, 19, 35] to CLIP‑based mesh highlighting [11, 8, 1]. All of these methods require either 2D mask supervision or large language–vision models, whereas our method learns part hierarchies directly from geometry without any auxiliary cues.
@@ 3 Method
In section˜3.1, we describe how decoder layers use part codebooks and cross-attention to define part–subpart relationships. In section˜3.2, we show how parts are geometrically grounded using 3D convex primitives with nested containment. Finally, section˜3.3 outlines the training objectives that combine reconstruction loss with regularizers to achieve self-supervised training.
@@ 4 Experiments
We validate our method through the task of part segmentation (section˜4.1), where we show our method can outperform all previous baselines in recovering consistent parts in all categories. We then show how our hierarchical part reconstruction achieves better reconstruction metrics in finer levels while higher levels remain semantically meaningful level of shape abstraction (section˜4.2). Finally, we ablate our design choices and analyze our training dynamics and recovered codebooks (section˜4.3).
@@ 4.1 Part segmentation – figures˜3 and 4
We train our method on $55$ categories from the ShapeNet-v2 dataset, following the train-test split used by Zhang et al. [50]. Our approach supports a unified training scheme, allowing all categories to be trained jointly. We further analyze this design in section˜4.3, where we compare it to category-specific training. For segmentation accuracy evaluation, since the ShapeNet [2] dataset does not provide segmentation labels, we leverage the fine-grained part annotations from the PartNet [27] dataset. This differs from previous methods, which evaluate segmentation accuracy on the ShapeNetPart dataset [47], containing only coarse segmentation ground-truth labels. The Door, Scissors, and Refrigerator categories appear only in PartNet and not in ShapeNet; hence for these, we use all instances for evaluation only. Each ground-truth point cloud for evaluation in PartNet dataset contains $10,000$ points.
@@
Our method significantly outperforms all baselines in part segmentation, both quantitatively and qualitatively as shown in Fig. 3 and 4. It produces consistent labeling of corresponding parts across instances within a category, whereas baselines often miss parts entirely or misclassify semantically similar regions.
@@ 4.2 Hierarchical reconstruction – figure˜5
We demonstrate that our model captures coherent high-level abstractions at early levels and progressively refines fine-grained details at deeper levels. Following the protocol of [50], we train and test on ShapeNet-v2 using the official splits. For evaluation, we compare reconstructed meshes—sampled to 100,000100{,}000 points each—against ground-truth point clouds, using symmetric Chamfer Distance and voxelized IoU at $128^{3}$ resolution. Our hierarchy employs $[6,16,24,36]$ parts. We compare against BAENet [7], RIMNet [30], and DAENet [5]. While BAENet and DAENet are single-level methods, RIMNet supports hierarchical reconstruction but only via a fixed binary tree, which yields coarse abstractions and loss of detail in later levels (figure˜5). By contrast, our model recovers flexible multi-branch hierarchies that balance abstraction and detail across levels.
4.3 Ablation study – tables˜1, 7, 6 and 8
@@
Figure 8: CD and IoU comparison with varying number of parts. As number of parts increases the reconstruction quality improves. $N_{l}$ is the number of parts at level 0. CD values are scaled by $0.1$.
We ablate our design choices by analyzing the effect of each loss, showing all loss components contribute to the quality of the decomposition. While balance and containment losses have little effect on quantitative metrics, removing them yields imbalanced and non–self-contained hierarchies, as seen in the qualitative results. While our method offers an effective, convenient, and generalized training on all object categories at once, we further analyze how a per-category training similar to previous works [7, 30, 5] affects our results. This shows that a per-category training can result in an improved specialized model for each category, in exchange for more time and compute. We provide a t-SNE analysis of the learned part embeddings, demonstrating that embeddings of similarly labeled parts cluster closely together. Finally, in figure˜8 we show that reconstruction quality improves as number of parts increases.
@@ 5 Conclusion
We introduce HiT, a self-supervised attention-based hierarchical neural field representation that learns general shape abstractions in a coarse-to-fine manner across diverse categories. Our core contribution is a novel cross-attention mechanism which enables dynamic discovery of parent-child relationships across hierarchy levels. To the best of our knowledge, our method is the first to enable general cross-category hierarchical 3D shape abstraction.
However, our current method is constrained by the requirement that the number of parts at each hierarchy level must be fixed. Such a constraint can sometimes lead to unnatural decompositions, particularly at finer levels of the hierarchy, limiting its effectiveness in downstream applications. Recent trends that leverage foundation models as semantic priors offer a promising way to address this limitation, as they enable semantically meaningful grouping of primitives based on high-level prior knowledge. A promising future direction is to make the number of convexes adaptive, for example by selecting them based on sparsity or reconstruction quality, so that the hierarchy remains compact and semantically meaningful.
Acknowledgements. We thank all the anonymous reviewers for their insightful comments. This research was supported by Adobe gift funds.
@@ Appendix A Supplementary
In the supplementary material, we present the hierarchical decomposition results of our method on the Objaverse dataset [9]. As shown in the examples below, our model dynamically allocates a different number of convexes depending on the input shape. Figures 9, 10, 11, 12, and 13 provide qualitative results of these decompositions. As can be seen, our model generates diverse decompositions tailored to the structural characteristics of each shape.